x[sup]2[/sup] - 6 = 3x + 9
1
Expert's answer
2011-01-27T08:33:29-0500
x2 - 6= 3x+9
x2 - 3x - 15= 0
D = 9 + 60 = 69
x1 = (3 + √69)/2 = 5.65
x2 = (3 - √69)/2 =-2.65